Use scientific notation to perform the following compatains
1. 80.000×5.000
2. 0.0006×8.000
1
Expert's answer
2020-09-17T09:28:45-0400

1. In scientific notation we have: "80 = 8\times 10^1" and "5 =5". Thus:


"8\times 10^1\cdot 5 = 8\cdot 5\times 10^1 = 40\times 10^1 = 4\times 10^2"

2. In scientific notation we have: "0.0006=6\times 10^{-4}" and "8 = 8". Thus:


"6\times 10^{-4}\cdot 8 = 6\cdot 8\times 10^{-4} = 48\times 10^{-4} = 4.8\times 10^{-3}"

Answer. 1) "4\times 10^2", 2) "4.8\times 10^{-3}" .